# Minimal Torsion Curves in Geometric Isogeny Classes

### v1

@@ 9.1. $\ell$ split in $K$
and this is attained by $E\in\mathcal{E}$ with CM by the maximal order in $K$. Thus $[\mathbb{Q}(j_{min}):\mathbb{Q}]=h_{K}$.
@@
That such an $E$ gives a point on $X_{1}(\ell^{n})$ in this degree follows from [4, Theorem 6.2] and $[\mathbb{Q}(j(E)):\mathbb{Q}]=h_{K}$ by Section 2.5; note that $\ell>3$ if $\Delta=-3,-4$ by the assumption that $\ell$ is split in $K$. It remains to show this is the least possible degree among all $E^{\prime}\in\mathcal{E}$. We may assume $\ell^{n}>2$. Since the endomorphism algebra is an isogeny invariant, any $E^{\prime}\in\mathcal{E}$ has CM by an order in $K$. Since we already have the least degree for a point with CM by the maximal order, we will henceforth assume $E^{\prime}$ has CM by an order in $K$ of conductor $\mathfrak{f}>1$. For any point $x=[E^{\prime},P^{\prime}]\in X_{1}(\ell^{n})$, by [3, Theorem 6.2] we have
@@
If $\deg(x)=h_{K}\cdot\frac{\ell^{n-1}(\ell-1)}{2}\cdot d<2\cdot h_{K}\cdot\frac{\ell^{n-1}(\ell-1)}{w_{K}}$ for some $d\in\mathbb{Z}^{+}$, it must be that $d=1$ and $w_{K}=2$. This implies $[\mathbb{Q}(j(E^{\prime})):\mathbb{Q}]=h_{K}$. The degree of this extension is equal to the class number of the order $\mathcal{O}$; see Equation 1 in Section 2.5. Since $w_{K}=2$, then $[\mathbb{Q}(j(E^{\prime})):\mathbb{Q}]=h_{K}$ implies $E^{\prime}$ has CM by an order in $K$ of conductor dividing 2. Since we have assumed $E^{\prime}$ has CM by an order of conductor $\mathfrak{f}>1$, we will suppose $E^{\prime}$ has CM by the order in $K$ of conductor 2. By Equation 1, this can happen only if 2 is split in $K$. But this contradicts [4, Theorem 6.2] if $\ell$ is odd and [4, Theorem 6.6] if $\ell=2$. ∎
@@ 9.2. $\ell$ inert in $K$
This is attained by $E\in\mathcal{E}$ with CM by an order in $K$ of conductor $\mathfrak{f}=\ell^{\lfloor{n/2}\rfloor}$. Moreover, if $j_{min}$ is the $j$-invariant of a minimal torsion curve of level $\ell^{n}$, then $[\mathbb{Q}(j_{min}):\mathbb{Q}]\rightarrow\infty$ as $n\rightarrow\infty$.
@@
Suppose $E$ has CM by the order in $K$ of conductor $\mathfrak{f}=\ell^{\lfloor{n/2}\rfloor}$. Note we can find such an $E\in\mathcal{E}$ by the first paragraph of §$\S 9$. Then by [4, Theorem 6.1, Theorem 6.6], the point $x\in X_{1}(\ell^{n})$ of least degree associated to such an $E$ has
@@
where $T(\mathcal{O},\ell^{n})$ is as defined in [4, Theorem 4.1] and $\epsilon=1$ if $\ell=2$, $n>1$ and $\epsilon=0$ otherwise. Evaluating $T(\mathcal{O},\ell^{n})$ via [4, Theorem 4.1] and $h(\mathcal{O})$ with Equation 1 shows $\deg(x)=\delta$.
Now we will justify that this is the least possible degree of a point on $X_{1}(\ell^{n})$ associated to $\mathcal{E}$. Suppose $E^{\prime}\in\mathcal{E}$ has CM by the order of conductor $\ell^{c}\mathfrak{f}^{\prime}$ in $\mathcal{O}_{K}$ where $\ell\nmid\mathfrak{f}^{\prime}$. By [4, Theorem 4.1, Theorem 6.1] and Equation 1, the least degree of a point on $X_{1}(\ell^{n})$ associated to $E^{\prime}$ is at least $\delta$, and this inequality is strict if $c<\frac{n-3}{2}$. Thus any minimal torsion curve of level $\ell^{n}$ must have $c\geq\frac{n-3}{2}$ and so $[\mathbb{Q}(j_{min}):\mathbb{Q}]\rightarrow\infty$ as $n\rightarrow\infty$. ∎
@@ 9.3. $\ell$ ramified in $K$
The least degree is attained by $E\in\mathcal{E}$ with CM by an order in $K$ of conductor $\mathfrak{f}=\ell^{\lfloor{n/2}\rfloor}$. Also, for any $j$-invariant $j_{min}$ of a minimal torsion curve of level $\ell^{n}$, one has $[\mathbb{Q}(j_{min}):\mathbb{Q}]\rightarrow\infty$ as $n\rightarrow\infty$
@@
Suppose $E$ has CM by an order in $K$ of conductor $\mathfrak{f}=\ell^{\lfloor{n/2}\rfloor}$; we can find such an $E\in\mathcal{E}$ by the first paragraph of §$\S 9$. Then by [4, Theorem 6.6] the least degree $x\in X_{1}(\ell^{n})$ associated to $E$ is
@@
where $T(\mathcal{O},\ell^{n})$ is as defined in [4, Theorem 4.1] and $\epsilon=1$ if $\operatorname{ord}_{2}(\Delta_{K})=2,\ell=2,n$ an odd integer greater than 1 and $\epsilon=0$ otherwise. Evaluating $T(\mathcal{O},\ell^{n})$ via [4, Theorem 4.1] and replacing $h(\mathcal{O})$ with the formula in Equation 1 of §$\S 2.5$ shows $\deg(x)=\delta$.
Now we will justify that this is the least possible degree of a point on $X_{1}(\ell^{n})$ associated to $\mathcal{E}$. Suppose $E^{\prime}\in\mathcal{E}$ has CM by the order of conductor $\ell^{c}\mathfrak{f}^{\prime}$ in $\mathcal{O}_{K}$ where $\ell\nmid\mathfrak{f}^{\prime}$. By [4, Theorem 4.1, Theorem 6.6] and Equation 1, the least degree of a point on $X_{1}(\ell^{n})$ associated to $E^{\prime}$ is at least $\delta$, and this inequality is strict if $c<\frac{n-3}{2}$. Thus any minimal torsion curve of level $\ell^{n}$ must have $c\geq\frac{n-3}{2}$, meaning $[\mathbb{Q}(j_{min}):\mathbb{Q}]\rightarrow\infty$ as $n\rightarrow\infty$. ∎

### v2

@@ 9.1. $\ell$ split in $K$
and this is attained by $E\in\mathcal{E}$ with CM by the maximal order in $K$. Thus $\mathbb{Q}(j_{min}):\mathbb{Q}]=h_{K}$.
@@
That such an $E$ gives a point on $X_{1}(\ell^{n})$ in this degree follows from [BC2, Theorem 6.2] and $\mathbb{Q}(j(E)):\mathbb{Q}]=h_{K}$ by Section 2.5; note that $\ell>3$ if $\Delta=-3,-4$ by the assumption that $\ell$ is split in $K$. It remains to show this is the least possible degree among all $E^{\prime}\in\mathcal{E}$. We may assume $\ell^{n}>2$. Since the endomorphism algebra is an isogeny invariant, any $E^{\prime}\in\mathcal{E}$ has CM by an order in $K$. Since we already have the least degree for a point with CM by the maximal order, we will henceforth assume $E^{\prime}$ has CM by an order in $K$ of conductor $\mathfrak{f}>1$. For any point $x=[E^{\prime},P^{\prime}]\in X_{1}(\ell^{n})$, by [BC1, Theorem 6.2] we have
@@
If $\deg(x)=h_{K}\cdot\frac{\ell^{n-1}(\ell-1)}{2}\cdot d<2\cdot h_{K}\cdot\frac{\ell^{n-1}(\ell-1)}{w_{K}}$ for some $d\in\mathbb{Z}^{+}$, it must be that $d=1$ and $w_{K}=2$. This implies $\mathbb{Q}(j(E^{\prime})):\mathbb{Q}]=h_{K}$. The degree of this extension is equal to the class number of the order $\mathcal{O}$; see Equation 1 in Section 2.5. Since $w_{K}=2$, then $\mathbb{Q}(j(E^{\prime})):\mathbb{Q}]=h_{K}$ implies $E^{\prime}$ has CM by an order in $K$ of conductor dividing 2. Since we have assumed $E^{\prime}$ has CM by an order of conductor $\mathfrak{f}>1$, we will suppose $E^{\prime}$ has CM by the order in $K$ of conductor 2. By Equation 1, this can happen only if 2 is split in $K$. But this contradicts [BC2, Theorem 6.2] if $\ell$ is odd and [BC2, Theorem 6.6] if $\ell=2$. ∎
@@ 9.2. $\ell$ inert in $K$
This is attained by $E\in\mathcal{E}$ with CM by an order in $K$ of conductor $\mathfrak{f}=\ell^{\lfloor{n/2}\rfloor}$. Moreover, if $j_{min}$ is the $j$-invariant of a minimal torsion curve of level $\ell^{n}$ and $n\geq 5$, then
$\mathbb{Q}(j_{min}):\mathbb{Q}]\geq h_{K}\cdot\frac{\ell^{(n-5)/2}}{3}(\ell+1).$
Thus $\mathbb{Q}(j_{min}):\mathbb{Q}]\rightarrow\infty$ as $n\rightarrow\infty$.
@@
Suppose $E$ has CM by the order in $K$ of conductor $\mathfrak{f}=\ell^{\lfloor{n/2}\rfloor}$. Note we can find such an $E\in\mathcal{E}$ by the first paragraph of $\S 9$. Then by [BC2, Theorem 6.1, Theorem 6.6], the point $x\in X_{1}(\ell^{n})$ of least degree associated to such an $E$ has
@@
where $T(\mathcal{O},\ell^{n})$ is as defined in [BC2, Theorem 4.1] and $\epsilon=1$ if $\ell=2$, $n>1$ and $\epsilon=0$ otherwise. Evaluating $T(\mathcal{O},\ell^{n})$ via [BC2, Theorem 4.1] and $h(\mathcal{O})$ with Equation 1 shows $\deg(x)=\delta$.
Now we will justify that this is the least possible degree of a point on $X_{1}(\ell^{n})$ associated to $\mathcal{E}$. Suppose $E^{\prime}\in\mathcal{E}$ has CM by the order of conductor $\ell^{c}\mathfrak{f}^{\prime}$ in $\mathcal{O}_{K}$ where $\ell\nmid\mathfrak{f}^{\prime}$. By [BC2, Theorem 4.1, Theorem 6.1] and Equation 1, the least degree of a point on $X_{1}(\ell^{n})$ associated to $E^{\prime}$ is at least $\delta$, and this inequality is strict if $c<\frac{n-3}{2}$. Thus, any minimal torsion curve of level $\ell^{n}$ must have $c\geq\frac{n-3}{2}$. If $n\geq 5$, then $\mathbb{Q}(j_{min}):\mathbb{Q}]\geq h_{K}\cdot\frac{\ell^{(n-5)/2}}{3}(\ell+1)$ by Equation 1 in Section 2.5. The conclusion follows. ∎
@@ 9.3. $\ell$ ramified in $K$
The least degree is attained by $E\in\mathcal{E}$ with CM by an order in $K$ of conductor $\mathfrak{f}=\ell^{\lfloor{n/2}\rfloor}$. Also, for any $j$-invariant $j_{min}$ of a minimal torsion curve of level $\ell^{n}$ and $n\geq 5$, one has
$\mathbb{Q}(j_{min}):\mathbb{Q}]\geq h_{K}\cdot\frac{\ell^{(n-3)/2}}{3}.$
Thus $\mathbb{Q}(j_{min}):\mathbb{Q}]\rightarrow\infty$ as $n\rightarrow\infty$.
@@
Suppose $E$ has CM by an order in $K$ of conductor $\mathfrak{f}=\ell^{\lfloor{n/2}\rfloor}$; we can find such an $E\in\mathcal{E}$ by the first paragraph of $\S 9$. Then by [BC2, Theorem 6.6] the least degree $x\in X_{1}(\ell^{n})$ associated to $E$ is
@@
where $T(\mathcal{O},\ell^{n})$ is as defined in [BC2, Theorem 4.1] and $\epsilon=1$ if $\operatorname{ord}_{2}(\Delta_{K})=2,\ell=2,n$ an odd integer greater than 1 and $\epsilon=0$ otherwise. Evaluating $T(\mathcal{O},\ell^{n})$ via [BC2, Theorem 4.1] and replacing $h(\mathcal{O})$ with the formula in Equation 1 of $\S 2.5$ shows $\deg(x)=\delta$.
Now we will justify that this is the least possible degree of a point on $X_{1}(\ell^{n})$ associated to $\mathcal{E}$. Suppose $E^{\prime}\in\mathcal{E}$ has CM by the order of conductor $\ell^{c}\mathfrak{f}^{\prime}$ in $\mathcal{O}_{K}$ where $\ell\nmid\mathfrak{f}^{\prime}$. By [BC2, Theorem 4.1, Theorem 6.6] and Equation 1, the least degree of a point on $X_{1}(\ell^{n})$ associated to $E^{\prime}$ is at least $\delta$, and this inequality is strict if $c<\frac{n-3}{2}$. Thus any minimal torsion curve of level $\ell^{n}$ must have $c\geq\frac{n-3}{2}$, meaning $\mathbb{Q}(j_{min}):\mathbb{Q}]\geq h_{K}\cdot\frac{\ell^{(n-3)/2}}{3}$ by Equation 1 in Section 2.5. The conclusion follows. ∎